Llong is 5 ft tall and is sanding in the light of a 15-ft lamppost. Her shadow is 4 ft long. If she walks 1 ft farther away from the lamppost, by how much will her shadow lengthen?
Answer by richwmiller(17219)

Here the height of the lamppost extra info. We don't use it.
5/4=x/5
5=4x/5
25=4x
25/4=x
It doesn't ask the length of her new shadow but how much is add to the old one.
25/4-16/4=9/4=2.25 feet
BTW The new shadow is 6.25 feet-4 =2.25 feet lengthened.